# ZEUS: Zero-shot Embeddings for Unsupervised Separation of Tabular Data

### v1

@@ 3.2 How effective is ZEUS for clustering?
To evaluate performance of the clustering methods, we employ a standard evaluation procedure, where clusters identified by models are expected to correspond to undisclosed ground-truth classes. We use the Adjusted Rand Index (ARI) for quantitative evaluation, and to improve readability, we scale the ARI values by a factor of 100, where 100 indicates perfect clustering and values near 0 represent random grouping. Table 1 presents a summary of ARI scores averaged over 5 random seeds and all datasets within each dataset group. Additionally, Table 2 displays the average rankings for each of the methods.
@@
KM GMM AE-KM AE-GMM DEC IDEC G-CEALS ZEUS
Real 55.54 48.49 51.43 53.56 55.93 54.57 40.37 57.43
Syn. Gauss. 89.90 76.93 81.26 81.40 89.36 82.57 62.84 89.03
Syn. Transf. 75.04 75.88 60.45 71.29 79.94 61.26 49.17 86.33
@@
KM GMM AE-KM AE-GMM DEC IDEC G-CEALS ZEUS
Real 3.90 4.57 4.72 4.29 3.96 4.31 6.65 3.60
Syn. Gauss. 2.60 3.35 5.40 5.45 3.10 5.50 7.70 2.90
Syn. Transf. 4.45 3.20 6.20 4.25 3.00 5.70 7.05 2.15
@@
Figure 3: Clustering time vs. input size.
Among the AE-based methods, DEC performs the best, while G-CEALS consistently demonstrates the weakest performance across all datasets. K-means and GMM, as representatives of the classical methods, perform reasonably well. In particular, k-means consistently lands among the best performing methods, which justifies its broad adoption among practitioners, despite being one of the most basic approaches. The detailed results including scores for individuals datasets corresponding to the averages in Tables 1 and 2 can be found in Appendix F.1
Finally, Figure 3 illustrates how the examined clustering methods scale with the increasing number of input data points. ZEUS maintains almost constant time while being only slightly slower than the basic k-means. It shows that the overhead from creating the representations by ZEUS is minimal. On the other hand, the remaining deep clustering algorithms require significantly more time and scale poorly with the increasing input size.
@@ 3.3 Are ZEUS’s predictions calibrated?
Table 3 reports the average Brier score computed over 5 seeds for all datasets in respective groups. For the analysis, we used the vanilla variant of ZEUS without regularization, which is paired with the GMM clustering since k-means does not provide soft assignments. The covariances were constrained to be identities as implied by Eq. 3. All competing baselines, except for k-means, provide probabilistic cluster assignments, making the Brier score calculation straightforward for them. For k-means, we used one-hot encoding to represent its assignments as probabilities.
@@
ZEUS demonstrates outstanding performance, achieving the best results for the Synthetic Transformed datasets while ranking second on both the OpenML and Synthetic Gaussian collections. The benefits of using ZEUS representations are especially visible in comparison against the vanilla GMM. Although the GMM achieves the second-best score on transformed data, its performance across the remaining datasets is merely modest. Among the remaining baselines, IDEC exhibits an interesting pattern: despite relatively weak clustering performance in Table 1, it achieves the top position on real datasets according to the Brier score and shows marked improvement on synthetic data. DEC displays the opposite tendency, with calibration results significantly inferior to its strong ARI performance. While G-CEALS remains generally weaker than other methods, its calibration performance consistently exceeds its clustering results presented in Table 1. The K-means algorithm, both in its standard implementation and when applied to autoencoder embeddings, yields impressive scores on Gaussian-categorical mixture datasets – likely due to the inherent separability of these data structures, which allows even simple one-hot probability estimates to produce well-calibrated predictions.
@@ 3.4 How helpful is regularization for ZEUS?
Our main model, which during pre-training incorporates both regularizers, exhibits a clear performance advantage on real datasets and consistently holds second place for the synthetic ones. Table 4 shows that a model pre-trained using only $\mathcal{L}_{prob}$ and its variant with $\mathcal{L}_{sep}$ included are insufficient for properly separating the transformer’s representations for the clustering task, as they both struggle with clustering the synthetic datasets encountered during pre-training. On the other hand, the results for $\mathcal{L}_{prob}+\mathcal{L}_{cp}$ imply a positive impact by the compact loss component, $\mathcal{L}_{cp}$, especially for synthetic data, which however does not translate to strong performance on the OpenML benchmark. We conclude that all three loss components are necessary for good performance, with $\mathcal{L}_{cp}$ being crucial for pre-training and $\mathcal{L}_{sep}$ important for real data.
@@ 4 Related work
A pioneering Deep Embedded Clustering (DEC) xie2016unsuperviseddeepembeddingclustering improves the target data representation by training the autoencoder and computing soft assignments in its latent space via Student’s $t$-distribution. In a concurrent work yang2017towards , the authors perform joint dimensionality reduction using AE and k-means clustering in the latent space of AE. Improved DEC (IDEC) guo_improved_2017 extends DEC by jointly optimizing reconstruction and clustering objectives, while spectral variations replace k-means steps with graph-based updates 8970928 . G-CEALS rabbani2025deep replaces the $t$–distribution assumption with multivariate Gaussian clusters. To eliminate the assumption of explicit distribution prior, DEPICT dizaji2017deepclusteringjointconvolutional attaches a softmax layer on top of an embedding network and train with cross-entropy loss. Finally, IDC svirsky2023interpretable predicts interpretable cluster assignments at the instance and cluster levels.
However, most of these deep learning models require careful hyperparameter tuning and early stopping which is unrealistic in the fully unsupervised setting due to the lack of labels 9506051 ; sadeghi2024deepclusteringselfsupervisionusing . Moreover, the optimization process has to be performed on each dataset, which is often time-consuming. Although multiple deep clustering approaches are currently in use, most of them are designed for texts or images znalezniak2023contrastive ; cai2023semantic ; li2024image and cannot be directly adapted to tabular data due to the lack of a dominant neural architecture for heterogeneous tabular inputs grinsztajn2022tree ; mcelfresh2023neural .
@@
Self-supervised learning (SSL) has been transformative in domains like vision and language chen2020simple ; grill2020bootstrap but has struggled to show similar success in tabular data yoon2020vime ; zhu2023xtab . Large diversity of data and lack of pre-defined correlation between features prevents from designing universal pretext tasks or augmentations as well as transferability between domains zhao2024comparison . In-context learning seems to be a promising direction in representation learning, which enables a single pre-trained model to adapt to new tasks in a zero-shot setting brown2020languagemodelsfewshotlearners . TabPFN treats small tabular datasets as contexts of (features, labels) and achieves state-of-the-art classification in one forward pass tabpfn . Although theoretical analyses reveal that transformers can implicitly implement algorithms such as gradient descent in context, their use is currently restricted to supervised problems garg2023transformerslearnincontextcase .
@@ 5 Conclusion
This research is part of the project No. 2022/45/P/ST6/02969 co-funded by the National Science Centre and the European Union Framework Programme for Research and Innovation Horizon 2020 under the Marie Skłodowska-Curie grant agreement No. 945339.
For the purpose of Open Access, the authors have applied a CC-BY public copyright licence to any Author Accepted Manuscript (AAM) version arising from this submission.
@@
The research of M. Śmieja and P. Marszałek was supported by the National Science Centre (Poland), grant no. 2023/50/E/ST6/00169.
The research of J. Tabor was supported by the National Science Centre (Poland), grant no. 2023/49/B/ST6/01137.
Witold Wydmański is supported by the Ministry of Science grant no. PN/01/0195/2022 and NCN Sonata BIS grant number 2020/38/E/NZ2/00598.
@@ Appendix B Details of synthetic data generation process
We use three types of probabilistic models to generate data from a category:
@@
    Gaussian. The continuous features of each cluster follow a multivariate Gaussian distribution with carefully constructed means and covariance matrices. Cluster means are placed with a minimum Wasserstein-2 distance constraint to ensure adequate separation between clusters and to prevent trivial overlap cases. The exact minimum distance value varies between 0.5 and 1.0 and is randomly selected for each component independently in order to promote data diversity. The covariance structure for each Gaussian is generated through eigendecomposition, with eigenvalues sampled from a predefined range of $[0.005,0.05]$ to control the shape and orientation of the clusters. Additionally, to guarantee the presence of both full-rank Gaussians in higher dimensions and degenerate ones in lower dimensions, extra conditions are introduced to narrow the range of eigenvalues. These constraints are activated with probabilities of 0.25 and 0.2, respectively. The process of constructing Gaussian mixture is incremental. Starting with a mixture containing $k$ components, the addition of a new cluster involves initially placing it at position 0 and then shifting it in a randomly chosen direction until the distance requirement is satisfied. This entire approach allows us to model a variety of cluster geometries, from spherical to highly elongated elliptical shapes.
@@
    Categorical. We also incorporate categorical features alongside continuous ones by sampling from categorical distributions that are biased toward certain categories for each cluster. To generate these varied categorical probability patterns, we use the Dirichlet distribution. The resulting categorical variables are then converted to one-hot encoding and combined with the continuous features, producing mixed-type datasets that more closely resemble real-world tabular data. The probability of including categorical features is controlled by a $categorical\_chance$ parameter, set to 0.3. Up to 3 categorical variables may be added, each having between 2 and $max\_categories$ possible values, defined as 5 .
@@
    NN transformed. To create more complex, non-linearly separable cluster structures, we apply transformations to the numerical features using random neural networks with 3 to 6 layers. These transformations map the original data through several non-linear operations while preserving cluster identity information, producing datasets with more challenging decision boundaries. Following the approach of Invertible Residual Networks [1], we constrain the spectral norm of each transformation layer to be less than 1. In addition, standardization is applied between residual layers to ensure more stable transformations. To help preserve cluster separability, we append one-hot vectors encoding the component identity of each data point to the numerical variables before passing them through the random neural network. After the transformation, these extra dimensions are removed using PCA, restoring the data to its original dimensionality. This comprehensive approach enables the simulation of intricate, non-linear data manifold structures often presented in real-world clustering problems without introducing degenerate configurations.
Afterwards, all features undergo standardization and scaling to ensure numerical stability during training. Continuous features are normalized to the range $[-1,1]$, while ensuring that the relative separation between clusters is preserved.
By training on this diverse collection of synthetic datasets-ranging from well-separated Gaussian clusters to complex, transformed manifolds with mixed feature types-our model learns to identify meaningful cluster structures across a wide spectrum of data distributions. This comprehensive prior fitting enables ZEUS to adapt to previously unseen datasets at inference time without additional training.

### v2

@@ 3.2 How effective is ZEUS for clustering?
To evaluate the performance of the clustering methods, we employ a standard evaluation procedure, where clusters identified by models are expected to correspond to undisclosed ground-truth classes. We use the Adjusted Rand Index (ARI) for quantitative evaluation, and to improve readability, we scale the ARI values by a factor of 100, where 100 indicates perfect clustering and values near 0 represent random grouping. Table 1 presents a summary of ARI scores averaged over 5 random seeds and all datasets within each dataset group. Additionally, Table 2 displays the average rankings for each of the methods.
@@
KM GMM AE-KM AE-GMM DEC IDEC IDC G-CEALS TabPFN SCARF ZEUS
Real 55.54 48.49 51.43 53.56 55.93 54.57 52.28 40.37 31.32 26.95 57.43
Syn. Gauss. 89.90 76.93 81.26 81.40 89.35 82.57 66.43 62.84 55.97 8.32 89.03
Syn. Transf. 75.04 75.88 60.45 71.29 79.94 61.26 66.78 49.17 15.66 2.48 86.33
@@
KM GMM AE-KM AE-GMM DEC IDEC IDC G-CEALS TabPFN SCARF ZEUS
Real 4.69 5.65 5.72 5.24 4.69 5.01 5.62 8.18 8.22 8.85 4.13
Syn. Gauss. 2.65 3.65 5.65 5.60 3.23 5.70 7.95 8.90 8.75 11.00 2.92
Syn. Transf. 4.80 3.50 6.85 4.50 3.20 6.35 6.05 7.80 9.75 11.00 2.20
Figure 3: Clustering time vs. input size.
@@
Among the AE-based methods, DEC performs the best, while G-CEALS consistently demonstrates the weakest performance across all datasets. K-means and GMM, as representatives of the classical methods, perform reasonably well. In particular, k-means consistently lands among the best-performing methods, which justifies its broad adoption among practitioners, despite being one of the most basic approaches. On the other hand, approaches such as SCARF or TabPFN are well-suited for tasks like classification, thanks to their strong performance in feature extraction. However, without additional regularization during training or specialized post-hoc fine-tuning (such as with DEC), effective separation of their representations is difficult. The detailed results, including scores for individual datasets corresponding to the averages in Tables 1 and 2, can be found in Appendix F.1
Finally, Figure˜3 illustrates how the examined clustering methods scale with the increasing number of input data points. ZEUS maintains almost constant time while being only slightly slower than the basic k-means. It shows that the overhead from creating the representations by ZEUS is minimal. On the other hand, the remaining deep clustering algorithms require significantly more time and scale poorly with the increasing input size.
@@ 3.3 Are ZEUS’s predictions calibrated?
Table 3 reports the average Brier score computed over 5 seeds for all datasets in respective groups. For the analysis, we used the vanilla variant of ZEUS, which is paired with the GMM clustering since k-means does not provide soft assignments. The covariances were constrained to be identities as implied by Eq. 3. All competing baselines, except for k-means, provide probabilistic cluster assignments, making the Brier score calculation straightforward for them. For k-means, we used one-hot encoding to represent its assignments as probabilities.
@@
ZEUS demonstrates outstanding performance, achieving the best results for the Synthetic Transformed datasets while ranking second on both the OpenML and Synthetic Gaussian collections. The benefits of using ZEUS representations are especially visible in comparison with the vanilla GMM. Although the GMM achieves the second-best score on transformed data, its performance across the remaining datasets is merely modest. Among the remaining baselines, IDEC exhibits an interesting pattern: despite relatively weak clustering performance in Table 1, it achieves the top position on real datasets according to the Brier score and shows marked improvement on synthetic data. DEC displays the opposite tendency, with calibration results significantly inferior to its strong ARI performance. While G-CEALS remains generally weaker than other methods, its calibration performance consistently exceeds its clustering results presented in Table 1. The K-means algorithm, both in its standard implementation and when applied to autoencoder embeddings, yields impressive scores on Gaussian-categorical mixture datasets – likely due to the inherent separability of these data structures, which allows even simple one-hot probability estimates to produce well-calibrated predictions.
@@ 3.4 How helpful is regularization for ZEUS?
Our main model, which during pre-training incorporates both regularizers, exhibits a clear performance advantage on real datasets and consistently holds second place for the synthetic ones. Table 4 shows that a model pre-trained using only $\mathcal{L}_{prob}$ and its variant with $\mathcal{L}_{sep}$ included are insufficient for properly separating the transformer’s representations for the clustering task, as they both struggle with clustering the synthetic datasets encountered during pre-training. On the other hand, the results for $\mathcal{L}_{prob}+\mathcal{L}_{cp}$ imply a positive impact by the compact loss component, $\mathcal{L}_{cp}$, especially for synthetic data, which, however, does not translate to strong performance on the OpenML benchmark. We conclude that all three loss components are necessary for good performance, with $\mathcal{L}_{cp}$ being crucial for pre-training and $\mathcal{L}_{sep}$ important for real data.
@@
3.5 What data-generating prior is optimal?
As mentioned in Section 2.2, constructing an appropriate data-generating prior $p(D)$ is crucial for maximizing performance and generalizability of ZEUS. To validate this claim, we conduct an ablation study evaluating four prior designs: Gaussian + Categorical, NN-transformed + Categorical, Gaussian + NN-transformed, and Gaussian + NN-transformed + Categorical (the standard ZEUS model). Each configuration name directly reflects the combination of priors used for pre-training (see Section B.2 in the Appendix for further details). For a fair comparison, we keep the default hyperparameter setup unchanged, except for the prior itself.
Table 5: ARI scores (averages; in rows) for ZEUS variants pre-trained on different priors (columns).
Gauss. + Cat. NN-transf. + Cat. Gauss. + NN-transf. Gauss. + NN-transf. + Cat.
Real 40.59 50.90 52.00 57.43
Syn. Gauss. 92.61 89.90 75.25 89.03
Syn. Transf. 73.34 87.04 71.29 86.33
The results of the study are summarized in Table 5. As expected, the Gaussian + Categorical model achieves top performance on the Synthetic Gaussian datasets, while the NN-transformed + Categorical model excels for the Synthetic Transformed data. Notably, the latter also performs well on the Gaussian sets, as its prior is built upon transformed Gaussian mixtures. In contrast, the Gaussian + NN-transformed model yields relatively poor average ARI scores across both synthetic benchmarks. This drop is primarily caused by a significant decline in ARI performance on the categorical datasets. Nonetheless, its average rank remains consistently competitive. Finally, for the OpenML datasets, the best performance is achieved when pre-training includes all three types of priors, underscoring the complementary importance of each prior in the training process.
@@ 4 Related work
A pioneering Deep Embedded Clustering (DEC) xie2016unsuperviseddeepembeddingclustering improves the target data representation by training the autoencoder and computing soft assignments in its latent space via Student’s $t$-distribution. In a concurrent work yang2017towards , the authors perform joint dimensionality reduction using AE and k-means clustering in the latent space. Then, improved DEC (IDEC) guo_improved_2017 extends DEC by jointly optimizing reconstruction and clustering objectives, while spectral variants replace k-means steps with graph-based updates 8970928 . G-CEALS rabbani2025deep replaces the tt–distribution assumption with multivariate Gaussian clusters. DEPICT dizaji2017deepclusteringjointconvolutional attaches a softmax layer on top of an embedding network and trains with cross-entropy loss to eliminate the assumption of an explicit distribution prior. Finally, IDC idc predicts interpretable cluster assignments at the instance and cluster levels.
Most of these deep learning models require careful hyperparameter tuning and early stopping which is unrealistic in the fully unsupervised setting due to the lack of labels 9506051 ; sadeghi2024deepclusteringselfsupervisionusing . Moreover, the optimization process has to be performed for each dataset, which is often time-consuming. Although multiple deep clustering approaches are currently in use, most of them are designed for texts or images znalezniak2023contrastive ; cai2023semantic ; li2024image and cannot be directly adapted for tabular data due to the lack of a dominant neural architecture for heterogeneous tabular inputs grinsztajn2022tree ; mcelfresh2023neural .
@@
Self-supervised learning (SSL) has been transformative in domains like vision and language chen2020simple ; grill2020bootstrap but has struggled to show similar success in tabular data yoon2020vime ; zhu2023xtab . A large diversity of data and a lack of pre-defined correlation between features hinder the design of universal pretext tasks or augmentations, as well as the transfer learning between domains zhao2024comparison . On the other hand, in-context and zero-shot learning brown2020languagemodelsfewshotlearners enable the use of a single pre-trained model for multiple tasks out of the box, without any additional tuning. In particular, TabPFN treats small tabular datasets as contexts consisting of features along with labels, and achieves state-of-the-art classification in a single forward pass tabpfn . Although theoretical analyses reveal that transformers can implicitly implement algorithms such as gradient descent in context, their use is currently restricted to supervised problems garg2023transformerslearnincontextcase .
@@ 5 Conclusion
We are grateful to the Reviewers for their effort and insightful comments on the paper.
This research is part of the project No. 2022/45/P/ST6/02969 co-funded by the National Science Centre and the European Union Framework Programme for Research and Innovation Horizon 2020 under the Marie Skłodowska-Curie grant agreement No. 945339. For the purpose of Open Access, the authors have applied a CC-BY public copyright licence to any Author Accepted Manuscript (AAM) version arising from this submission.
@@
The research of P. Marszałek and M. Śmieja was supported by the National Science Centre (Poland), grant no. 2023/50/E/ST6/00169. The research of Witold Wydmański was supported by the Ministry of Science grant no. PN/01/0195/2022 and NCN Sonata BIS grant number 2020/38/E/NZ2/00598. The research of J. Tabor was supported by the National Science Centre (Poland), grant no. 2023/49/B/ST6/01137. Some experiments were performed on servers purchased with funds from the flagship project entitled “Artificial Intelligence Computing Center Core Facility” from the DigiWorld Priority Research Area within the Excellence Initiative – Research University program at Jagiellonian University in Krakow. We also gratefully acknowledge Polish high-performance computing infrastructure PLGrid (HPC Center: ACK Cyfronet AGH) for providing computer facilities and support within computational grant no. PLG/2024/017893.
@@ Appendix B Details of synthetic data generation process
B.1 Theory for synthetic-to-real data generalization
Theorem 1 (Univeral Approximation Theorem for Mixture Distributions).
Let $P=\sum_{i=1}^{k}\pi_{i}P_{i}$ be a mixture of distributions on $\mathbb{R}^{d}$, where each $P_{i}$ is a probability distribution, and $\pi_{i}>0$, $\sum_{i=1}^{k}\pi_{i}=1$.
Then there exists a mixture of Gaussians $Q=\sum_{i=1}^{k}\pi_{i}\mathcal{N}(\mu_{i},\Sigma_{i})$ and a neural network $F$ such that we can approximate $P$ with arbitrarily small error by the pushforward measure $F_{\#}Q$. Additionally, we can select $F$ so that $F_{\#}N(\mu_{i},\Sigma_{i})$ approximates $P_{i}$ with arbitrarily small error.
Proof.
We proceed in several steps.
Step 1: Approximation of individual components.
Making use of the universal approximation theorem, for each component $P_{i}$ there exists a neural network $F_{i}$ which transforms an arbitrary Gaussian $Q_{i}=\mathcal{N}(\mu_{i},\Sigma_{i})$ into $P_{i}$, i.e. $F_{i\#}Q_{i}\approx P_{i}$.
Step 2: Constructing a unified neural network.
We now define a single neural network $F$ that can represent all the $F_{i}$ networks. Define $F:\mathbb{R}^{d}\times\{1,\dots,k\}\to\mathbb{R}^{d}$ such that $F(z,i)\approx F_{i}(z)$. This can be implemented by encoding the index $i$ as a one-hot vector or embedding and concatenating it to $z$, enabling $F$ to learn the behavior of each $F_{i}$ conditioned on the index $i$. The network $F$ can be extended to the domain $\mathbb{R}^{d}\times\mathbb{R}$ in an arbitrary way.
Step 3: Approximation of the full mixture.
Since each $F_{\#}\mathcal{N}(\mu_{i},\Sigma_{i})$ approximates $P_{i}$, the mixture $F_{\#}Q=\sum_{i=1}^{k}\pi_{i}F_{\#}\mathcal{N}(\mu_{i},\Sigma_{i})$ approximates $P=\sum_{i=1}^{k}\pi_{i}P_{i}$.
Thus, we have constructed a mixture of Gaussians $Q$ and a neural network $F$ such that $F_{\#}Q$ approximates $P$, and each component $F_{\#}\mathcal{N}(\mu_{i},\Sigma_{i})$ approximates $P_{i}$ arbitrarily well. This completes the proof. ∎
B.2 Insights into data-generating priors
We use three types of probabilistic procedures to generate the prior data:
@@ B.2 Insights into data-generating priors
    Gaussian. The goal is to create a continuous data type where each cluster follows a multivariate Gaussian distribution with carefully designed means and covariance matrices. This construction process is incremental. Starting with a mixture containing $k$ components, the addition of a new cluster involves initially placing it at position 0 and then shifting it in a randomly chosen direction. The covariance structure for each Gaussian is generated through eigendecomposition, with eigenvalues sampled from a predefined range of $[0.005,0.05]$ to control the shape and orientation of the clusters. Additionally, to guarantee the presence of both full-rank Gaussians in higher dimensions and degenerate ones in lower dimensions, extra conditions are introduced to narrow the range of eigenvalues. These constraints are activated with probabilities of 0.25 and 0.2, respectively. To ensure adequate separation between clusters and to prevent trivial overlap, we apply a Wasserstein-2 distance constraint, requiring that the means $\mu_{i}$ and $\mu_{j}$ of any two clusters, along with their corresponding covariance matrices $\Sigma_{i}$ and $\Sigma_{j}$, are separated by at least a threshold $T$, as follows:
    $\|\mu_{i}-\mu_{j}\|^{2}_{2}+tr(\Sigma_{i}+\Sigma_{j}-2(\Sigma_{i}^{\frac{1}{2}}\Sigma_{j}\Sigma_{i}^{\frac{1}{2}})^{\frac{1}{2}})\geq T$ (7)
    The exact minimum distance value $T$ varies between 0.5 and 1.0 and is randomly selected for each component independently in order to promote data diversity. This entire approach allows us to model a variety of cluster geometries, from spherical to highly elongated elliptical shapes.
@@
    Categorical. To enrich the generated data, we incorporate categorical features alongside continuous ones by sampling from categorical distributions that are biased toward certain categories for each cluster. To produce these varied categorical probability patterns, we use the Dirichlet distribution. The probability of including categorical features is controlled by a categorical_chancecategorical\_chance parameter, set to 0.3. Up to 3 categorical variables may be added, each having between 2 and max_categoriesmax\_categories possible values, defined as 5. The resulting categorical variables are then converted to one-hot encoding and combined with the continuous features, producing mixed-type datasets that more closely resemble real-world tabular data.
@@
    NN transformed. To create more complex, non-linearly separable cluster structures, we apply transformations to the numerical features using random neural networks with 3 to 6 layers. These transformations map the original data through several non-linear operations while preserving cluster identity information, producing datasets with more challenging decision boundaries. Following the approach of Invertible Residual Networks [2], we constrain the spectral norm of each transformation layer to be less than 1. In addition, standardization is applied between residual layers to ensure more stable transformations. To help preserve cluster separability, we append one-hot vectors indicating the component identity to the numerical variables before passing them through the random neural network. After the transformation, these extra dimensions are removed using the PCA algorithm, restoring the data to its original dimensionality. This comprehensive approach enables the simulation of intricate, non-linear data manifold structures often presented in real-world clustering problems without introducing degenerate configurations.
All features undergo standardization and scaling to ensure numerical stability during training. Continuous features are normalized to the range [1,1][-1,1], while ensuring that the relative separation between clusters is preserved.
By training on this diverse collection of synthetic datasets-ranging from well-separated Gaussian clusters to complex, transformed manifolds with mixed feature types ZEUS learns to identify meaningful cluster structures across a wide spectrum of data distributions. This enables it to adapt to previously unseen datasets at inference time without additional training.